# Topological Interface States and Nonlinear Thermoelectric Performance in Armchair Graphene Nanoribbon Heterostructures

### v1

@@ III.4 Nonlinear thermoelectric power of TDQD
Here, $E_{j}$ is the spin-independent energy level of dot $j$, and $U_{j}=U_{L(R)}=U_{0}$ and $U_{j,\ell}=U_{LR}=U_{1}$ denote the intra-dot and inter-dot Coulomb interactions, respectively. The operator $n_{j,\sigma}=c^{\dagger}_{j,\sigma}c_{j,\sigma}$ is the number operator, and $t_{x}$ is the interdot hopping amplitude. Because the interface-state energy levels of 15-13-15 AGNRHs lie deep within the bulk band gap, thermal noise effects such as phonon-assisted excitations can be safely neglected in Eq. (3).
The Green-function technique provides a powerful framework for calculating tunneling currents in strongly correlated nanostructures under nonequilibrium conditions [DavidK1, 69Kuo6, 72]. Using the equation-of-motion method, the tunneling current from the left (right) electrode through the TDQD with electron-electron interactions is given by
@@
The Fermi-Dirac distribution of electrode $\alpha$ is defined as $f_{\alpha}(\varepsilon)=1/({\exp[(\varepsilon-\mu_{\alpha})/k_{B}T_{\alpha}]+1}$),with chemical potentials $\mu_{L(R)}=\mu\pm eV_{a}/2$ under an applied bias of $V_{a}/2$ and $V_{a}/2$. For convenience, we set the Fermi energy $\mu=0$. The temperature bias is defined as $\Delta T=T_{L}-T_{R}$. A closed-form expression for ${\cal T}_{LR}(\varepsilon)$ is given in Ref. [Kuo6, 72]:
@@
Here, $N_{\ell,\sigma}$ is the single-particle occupation at site $\ell$. Two-particle and three-particle correlation functions are also included, and the full set is solved self-consistently so that $\sum_{m}C_{m}=1$, ensuring probability conservation. The reverse-bias current is obtained by exchanging the indices of ${\cal T}_{LR}(\varepsilon)$. Note that the transmission formula is valid only for temperatures above the Kondo temperature [MadhavanV, 73GoldhaberG, 75].
Although many-body effects in voltage-driven transport have been widely explored, their impact on electrical power generation in nanostructures driven by a temperature bias–both in quasi-equilibrium and far-from-equilibrium regimes–remains insufficiently understood [ChenG, 76SobrinoN, 97]. Here, we investigate these many-body effects for thermoelectric generators formed by the topological states (TSs) of AGNRHs. To reflect realistic experimental conditions, the left and right TQD energy levels are modeled as $E_{L}=-(\eta_{1}eV_{L,g}+(1-\eta_{1})eV_{R,g})+\eta_{L}~eV_{a}$ and $E_{R}=-((1-\eta_{2})eV_{L,g}+\eta_{2}V_{R,g})+\eta_{R}~eV_{a}$, where $\eta_{1}$ and $\eta_{2}$ represent the effects of intra- and interdot Coulomb interactions [DasSarma, 98] and $\eta_{L(R)}$ ($|\eta_{L(R)}|<1/2$) describe the orbital shifts induced by $V_{a}$ (see Fig. (3)). We take $U_{0}=90$ meV and $U_{1}=25$ meV, consistent with their weak dependence on AGNR lengths [Kuo5, 56]. In this study, we use $\eta_{1}=\eta_{2}=0.8$ and a small orbital offset $\eta_{L}=-\eta_{R}=0.1$. The tunneling rates $\Gamma_{e,L}$ and $\Gamma_{e,R}$ and the interdot hopping $t_{x}$ are treated as tunable parameters controlled by the 13- and 15-AGNR segment lengths.
To analyze nonlinear temperature-driven power generation, we define the optimal electrical power as $\Omega_{op}=-J_{op}(V_{op},\Delta T)\times V_{op}(\Delta T)$, where $J_{op}(V_{op},\Delta T)$ and $V_{op}(\Delta T)$ denote the optimized thermal current and thermal voltage, respectively [JiangJH, 85,FastJ, 94]. The thermal voltage $V_{op}(\Delta T)$ is obtained by maximizing the electrical power $\Omega=-J_{L(R)}(V_{a},\Delta T)\times V_{a}$ with respect to $V_{a}$ [FastJ, 94]. This optimization is highly nontrivial in the Coulomb blockade regime under nonequilibrium conditions; thus, many theoretical studies avoid correlated transport and power optimization altogether [AzemaJ, 83SanchezR, 95]. The electrode temperatures are expressed as $T_{L}=T_{0}\pm\Delta T/2$ and $T_{R}=T_{0}\mp\Delta T/2$, where $T_{0}=(T_{L}+T_{R})/2$ is the average temperature.
@@ 2. Comparison between noninteracting and interacting cases
Many previous studies predicted the power output of nanoscale thermoelectric generators without considering electron-electron Coulomb interactions [Hershfield, 80RaymondJH, 96]. To clarify the role of Coulomb interactions and temperature bias on $V_{op}$ and $J_{op}$ at room temperature, Fig. 8 shows $V_{op}$, $J_{op}$, and $\Omega_{op}$ as functions of gate voltage $V_{g}$ and temperature bias $\Delta T$ at $T_{0}=288$ K. $V_{g}$ is varied from 100-100 mV to 0, and $\Delta T$ from 0 to 240 K. Fig. 8(a)-8(c) correspond to the noninteracting case, and Fig. 8(d)-8(f) to the interacting case. As seen in Fig. 8(a) and Fig. 8(d), $V_{op}$ differs only slightly, mainly near $V_{g}\geq-10$ mV: without interactions, $V_{op}(V_{g}=0)=0$ for all $\Delta T$ due to electron-hole symmetry, whereas interactions lift this symmetry, giving small but finite $V_{op}$ at $V_{g}=0$. In contrast, $J_{op}$ shows large differences: the noninteracting maximum $J_{op,\max}=0.566$ at $V_{g}=-25$ mV and $\Delta T=240$ K [Fig.8 (b)] is significantly overestimated compared with the interacting value $J_{op,\max}=0.418$ at $V_{g}=-32$ mV [Fig. 8(e)]. In the (0,0) configuration, the Coulomb blockade suppresses $J_{op,\max}$ but has little effect on $V_{op,\max}$. The behavior of $V_{op}$ resembles that of the Seebeck coefficient $S$, which is largely insensitive to electron Coulomb interactions [Kuo9, 78].
@@
To further clarify the relationship between $\Omega_{op}$ and $\Delta T$ shown in Fig. 8(f), Figs. 9(a)-9(c) present $S_{op}=V_{op}/\Delta T$, $J_{op}$, and $\Omega_{op}$ as functions of $\Delta T$ for $V_{g}=-32$, 46-46, and 60-60 mV. For $V_{g}=-60$ mV, $\Omega_{op}$ reaches its maximum value at $k_{B}\Delta T=20$ meV. As shown in Fig. 9(a), $S_{op}$ is only weakly dependent on $\Delta T$ for $V_{g}=-32$ mV, which also explains why $V_{op}$ exhibits a linear dependence on $\Delta T$ for $V_{g}>-24$ mV in Fig. 8(d). When the TDQD energy levels lie far from the Fermi energy ($V_{g}=-60$ mV), $S_{op}$ remains linear in $\Delta T$, and this linearity persists up to $k_{B}\Delta T=20$ meV. It is noteworthy that $k_{B}\Delta T/(k_{B}T_{0})\geq\frac{10}{24}\geq\frac{5}{12}$ is not a perturbative parameter, yet the nonlinear correction to $V_{op}$ is restricted to a quadratic term, with no observable cubic or higher-order contributions which is an unexpected behavior. This finding is consistent with the recent analytical results reported in Ref. [RaymondJH, 96], where a single noninteracting quantum dot with one energy level was considered. In Fig. 9(b), we find that $J_{op}$ remains a linear function of $\Delta T$ even at $V_{g}=-60$ mV. Based on the trends identified in Fig. 9(a), $\Omega_{op}=-J_{op}V_{op}$ for $V_{g}=-46$ and 60-60 mV, with $\Omega_{op,\max}$ deviating from the expected $(\Delta T)^{2}$ scaling due to nonlinearities in $V_{op}$.
@@ IV Conclusion
Using the calculated transmission coefficient ${\cal T}_{GNR}(\varepsilon)$ for $N_{O}$-$N_{C}-AGNR$$N_{O}$-AGNR segments with zigzag edge structures coupled to electrodes, we demonstrated that IFs act as a topological double quantum dot (TDQD) when the IFs are formed by the ESs of the central AGNR segment. By employing an Anderson model incorporating effective intradot and interdot Coulomb interactions, we derived an analytical expression for the tunneling current through the TDQD via the transmission coefficient ${\cal T}_{LR}(\varepsilon)$. The thermoelectric performance of AGNRH-based TDQDs was further analyzed in the context of their potential application as graphene-nanoribbon power generators [PerrinMM, 103].
@@ Appendix A Interface States of Wide-AGNR/Narrow-AGNR/Wide-AGNR AGNRH Segments
For N-AGNR/(N2)(N-2)-AGNR/N-AGNR AGNRH segments, the width difference $\Delta N=(N_{out}-N_{cen})/2$ equals $\Delta N=1$, where $N_{out}$ and $N_{cen}$ denote the atom widths of the outer and central AGNR segments, respectively. In Fig. 4, our analysis focused on the case $\Delta N=1$. In this appendix, we examine how increasing $\Delta N$ affects both the number and chirality of interface states (IFs) in AGNRHs.
Figures A.1(a-d) display the energy spectra of four AGNRH segments as functions of $t_{es}$: (a) $27_{4}-25_{10}-27_{4}$, (b) $27_{4}-21_{10}-27_{4}$, (c) $27_{4}-13_{10}-27_{4}$, and (d) $27_{4}-9_{10}-27_{4}$. These correspond to $\Delta N=1,3,7$, and $9$, respectively, with a fixed outer width of $N_{out}=27$. In all cases, the outer AGNRs have an $R_{1}$-type unit cell, whereas the central AGNR adopts an $R_{2}$-type unit cell.
In Fig. A.1(a), one interface state of sublattice-A (sublattice-B) arises at the 27-AGNR/25-AGNR (25-AGNR/27-AGNR) junction. The resulting pair of IF levels, $\Sigma_{IF,c}$ and $\Sigma_{IF,v}$, originates from their mutual coupling. Compared with the $27_{4}-25_{6}-27_{4}$ case in Fig. 4(d), the energy separation between $\Sigma_{IF,c}$ and $\Sigma_{IF,v}$ in Fig. A.1(a) is smaller due to the longer $25_{10}$ central segment.
In Fig. A.1(b), eight in-gap states appear, all attributable to the end states of the 27212727-21-27 AGNRH system. Since both 27-AGNR and 21-AGNR segments have $N_{A(B)}=4$, the number of IFs is zero, consistent with the rule that IFs are determined by the difference in ES counts.
In Fig. A.1(c), a single interface state with sublattice-B(A) character is present at each 27-AGNR/13-AGNR (13-AGNR/27-AGNR) junction because the 13-AGNR segment has $N_{A(B)}=3$. These IFs originate from the ESs of the outer AGNRs; consequently, their mutual overlap is extremely small, whereas the IFs strongly couple to the terminal states of the 27-13-27 AGNRH segment. Their evolution with respect to $t_{es}$ is therefore markedly different from IFs originating from the central AGNR segment.
In Fig. A.1(d), two IFs appear at the 27-AGNR/9-AGNR (9-AGNR/27-AGNR) junctions because the 9-AGNR segment possesses $N_{A(B)}=2$.
@@
Figure A.3 shows the probability densities of representative IF levels specifically, the pair $\Sigma_{IF,c2,a}=102.2335$ meV and $\Sigma_{IF,c2,b}=102.2232$ meV, the level $\Sigma_{IF,c1}=15.4212$ meV, and the terminal-state level $\Sigma_{A,c}=0.539$ meV corresponding to Fig. A.2(a). In Figs. A.3(a) and A.3(b), the spatial distributions of $\Sigma_{IF,c2,a}$ and $\Sigma_{IF,c2,b}$ are nearly indistinguishable; furthermore, their amplitudes inside the central 9-AGNR segment are extremely weak, indicating that transport through $\Sigma_{IF,c2}$ is strongly suppressed.
In contrast, $\Sigma_{IF,c1}$ exhibits a pronounced probability density within the central 9-AGNR segment, making it a more favorable transport channel. As shown in Fig. A.3(c), $\Sigma_{IF,c1}$ effectively behaves as a set of four topological dots (TDs). Figure A.3(d) demonstrates that the state $\Sigma_{A,c}$ is primarily associated with the terminal states of the 27-9-27 AGNRH structure.

### v2

@@ III.4 Nonlinear thermoelectric power of TDQD
Here, $E_{j}$ is the spin-independent energy level of dot $j$, and $U_{j}=U_{L(R)}=U_{0}$ and $U_{j,\ell}=U_{LR}=U_{1}$ denote the intra-dot and inter-dot Coulomb interactions, respectively. The operator $n_{j,\sigma}=c^{\dagger}_{j,\sigma}c_{j,\sigma}$ is the number operator, and $t_{x}$ is the interdot hopping amplitude. For 15-13-15 AGNRH segments, both end states and interface states emerge within the middle gap and are well separated from the conduction and valence subbands by a band gap of approximately 1 eV [Fig. 3(b)]. This large separation suppresses optical-phonon assisted transport due to the phonon bottleneck effect [FangJW, 69]; therefore, thermal noise effects arising from bound-to-continuum states can be safely neglected in Eq. (3). Noting that the phonon mean free path in graphene nanostructures can be reduced to 10 nm [ZuevYM, 70-HsiehYP, 76]. Since the lengths of the AGNRH segments considered in this work are smaller than this scale, electron-acoustic phonon scattering can be safely ignored.
The Green-function technique provides a powerful framework for calculating tunneling currents in strongly correlated nanostructures under nonequilibrium conditions [DavidK1, 77Kuo6, 80]. Using the equation-of-motion method, the tunneling current from the left (right) electrode through the TDQD with electron-electron interactions is given by
@@
The Fermi-Dirac distribution of electrode $\alpha$ is defined as $f_{\alpha}(\varepsilon)=1/({\exp[(\varepsilon-\mu_{\alpha})/k_{B}T_{\alpha}]+1}$),with chemical potentials $\mu_{L(R)}=\mu\pm eV_{a}/2$ under an applied bias of $V_{a}/2$ and $V_{a}/2$. For convenience, we set the Fermi energy $\mu=0$. The temperature bias is defined as $\Delta T=T_{L}-T_{R}$. A closed-form expression for ${\cal T}_{LR}(\varepsilon)$ is given in Ref. [Kuo6, 80]:
@@
Here, $N_{\ell,\sigma}$ is the single-particle occupation at site $\ell$. Two-particle and three-particle correlation functions are also included, and the full set is solved self-consistently so that $\sum_{m}C_{m}=1$, ensuring probability conservation. The reverse-bias current is obtained by exchanging the indices of ${\cal T}_{LR}(\varepsilon)$. Note that the transmission formula is valid only for temperatures above the Kondo temperature [MadhavanV, 81GoldhaberG, 83].
Although many-body effects in voltage-driven transport have been widely explored, their impact on electrical power generation in nanostructures driven by a temperature bias–both in quasi-equilibrium and far-from-equilibrium regimes–remains insufficiently understood [ChenG, 84SobrinoN, 106]. Here, we investigate these many-body effects for thermoelectric generators formed by the topological states (TSs) of AGNRHs. In realistic experimental conditions, the left (right) gate voltage not only tunes the energy level of the left (right) TQD but also affects the right (left) TQD; therefore, the left and right TQD energy levels are modulated as $E_{L}=-(\eta_{1}eV_{L,g}+(1-\eta_{1})eV_{R,g})+\eta_{L}~eV_{a}$ and $E_{R}=-((1-\eta_{2})eV_{L,g}+\eta_{2}V_{R,g})+\eta_{R}~eV_{a}$, where $\eta_{1}$ and $\eta_{2}$ represent the effects of intra- and interdot Coulomb interactions [DasSarma, 107] and $\eta_{L(R)}$ ($|\eta_{L(R)}|<1/2$) describe the orbital shifts induced by $V_{a}$ (see Fig. (3)). We take $U_{0}=90$ meV and $U_{1}=25$ meV, consistent with their weak dependence on AGNR lengths [Kuo5, 56]. In this study, we use $\eta_{1}=\eta_{2}=0.8$ and a small orbital offset $\eta_{L}=-\eta_{R}=0.1$. The tunneling rates $\Gamma_{e,L}$ and $\Gamma_{e,R}$ and the interdot hopping $t_{x}$ are treated as tunable parameters controlled by the 13- and 15-AGNR segment lengths.
To analyze nonlinear temperature-driven power generation, we define the optimal electrical power as $\Omega_{op}=-J_{op}(V_{op},\Delta T)\times V_{op}(\Delta T)$, where $J_{op}(V_{op},\Delta T)$ and $V_{op}(\Delta T)$ denote the optimized thermal current and thermal voltage, respectively [JiangJH, 93,FastJ, 103]. The thermal voltage $V_{op}(\Delta T)$ is obtained by maximizing the electrical power $\Omega=-J_{L(R)}(V_{a},\Delta T)\times V_{a}$ with respect to $V_{a}$ [FastJ, 103]. This optimization is highly nontrivial in the Coulomb blockade regime under nonequilibrium conditions; thus, many theoretical studies avoid correlated transport and power optimization altogether [AzemaJ, 91SanchezR, 104]. The electrode temperatures are expressed as $T_{L}=T_{0}\pm\Delta T/2$ and $T_{R}=T_{0}\mp\Delta T/2$, where $T_{0}=(T_{L}+T_{R})/2$ is the average temperature.
@@ 2. Comparison between noninteracting and interacting cases
Many previous studies predicted the power output of nanoscale thermoelectric generators without considering electron-electron Coulomb interactions [Hershfield, 88RaymondJH, 105]. To clarify the role of Coulomb interactions and temperature bias on $V_{op}$ and $J_{op}$ at room temperature, Fig. 8 shows $V_{op}$, $J_{op}$, and $\Omega_{op}$ as functions of gate voltage $V_{g}$ and temperature bias $\Delta T$ at $T_{0}=288$ K. $V_{g}$ is varied from 100-100 mV to 0, and $\Delta T$ from 0 to 240 K. Fig. 8(a)-8(c) correspond to the noninteracting case, and Fig. 8(d)-8(f) to the interacting case. As seen in Fig. 8(a) and Fig. 8(d), $V_{op}$ differs only slightly, mainly near $V_{g}\geq-10$ mV: without interactions, $V_{op}(V_{g}=0)=0$ for all $\Delta T$ due to electron-hole symmetry, whereas interactions lift this symmetry, giving small but finite $V_{op}$ at $V_{g}=0$. In contrast, $J_{op}$ shows large differences: the noninteracting maximum $J_{op,\max}=0.566$ at $V_{g}=-25$ mV and $\Delta T=240$ K [Fig.8 (b)] is significantly overestimated compared with the interacting value $J_{op,\max}=0.418$ at $V_{g}=-32$ mV [Fig. 8(e)]. In the (0,0) configuration, the Coulomb blockade suppresses $J_{op,\max}$ but has little effect on $V_{op,\max}$. The behavior of $V_{op}$ resembles that of the Seebeck coefficient $S$, which is largely insensitive to electron Coulomb interactions [Kuo9, 86].
@@
[…]
@@ IV Conclusion
Using the calculated transmission coefficient ${\cal T}_{GNR}(\varepsilon)$ for $N_{out}$-$N_{cen}-AGNR$$N_{out}$-AGNR segments with zigzag edge structures coupled to electrodes, we demonstrated that IFs act as a topological double quantum dot (TDQD) when the IFs are formed by the ESs of the central AGNR segment. By employing an Anderson model incorporating effective intradot and interdot Coulomb interactions, we derived an analytical expression for the tunneling current through the TDQD via the transmission coefficient ${\cal T}_{LR}(\varepsilon)$. The thermoelectric performance of AGNRH-based TDQDs was further analyzed in the context of their potential application as graphene-nanoribbon power generators [PerrinMM, 113].
@@ Appendix A Interface States of Wide-AGNR/Narrow-AGNR/Wide-AGNR AGNRH Segments
For N-AGNR/(N2)(N-2)-AGNR/N-AGNR AGNRH segments, the width difference $\Delta N=(N_{out}-N_{cen})/2$ equals $\Delta N=1$, where $N_{out}$ and $N_{cen}$ denote the atom widths of the outer and central AGNR segments, respectively. In Fig. 4, our analysis focused on the case $\Delta N=1$. In this appendix, we examine how increasing $\Delta N$ affects both the number and chirality of interface states (IFs) in AGNRHs. Figures A.1(a-d) display the energy spectra of four AGNRH segments as functions of $t_{es}$: (a) $27_{4}-25_{10}-27_{4}$, (b) $27_{4}-21_{10}-27_{4}$, (c) $27_{4}-13_{10}-27_{4}$, and (d) $27_{4}-9_{10}-27_{4}$. These correspond to $\Delta N=1,3,7$, and $9$, respectively, with a fixed outer width of $N_{out}=27$. In all cases, the outer AGNRs have an $R_{1}$-type unit cell, whereas the central AGNR adopts an $R_{2}$-type unit cell.
In Fig. A.1(a), one interface state of sublattice-A (sublattice-B) arises at the 27-AGNR/25-AGNR (25-AGNR/27-AGNR) junction. The resulting pair of IF levels, $\Sigma_{IF,c}$ and $\Sigma_{IF,v}$, originates from their mutual coupling. Compared with the $27_{4}-25_{6}-27_{4}$ case in Fig. 4(d), the energy separation between $\Sigma_{IF,c}$ and $\Sigma_{IF,v}$ in Fig. A.1(a) is smaller due to the longer $25_{10}$ central segment. In Fig. A.1(b), eight in-gap states appear, all attributable to the end states of the 27212727-21-27 AGNRH system. Since both 27-AGNR and 21-AGNR segments have $N_{A(B)}=4$, the number of IFs is zero, consistent with the rule that IFs are determined by the difference in ES counts. In Fig. A.1(c), a single interface state with sublattice-B(A) character is present at each 27-AGNR/13-AGNR (13-AGNR/27-AGNR) junction because the 13-AGNR segment has $N_{A(B)}=3$. These IFs originate from the ESs of the outer AGNRs; consequently, their mutual overlap is extremely small, whereas the IFs strongly couple to the terminal states of the 27-13-27 AGNRH segment. Their evolution with respect to $t_{es}$ is therefore markedly different from IFs originating from the central AGNR segment. In Fig. A.1(d), two IFs appear at the 27-AGNR/9-AGNR (9-AGNR/27-AGNR) junctions because the 9-AGNR segment possesses $N_{A(B)}=2$.
@@
Figure A.3 shows the probability densities of representative IF levels specifically, the pair $\Sigma_{IF,c2,a}=102.2335$ meV and $\Sigma_{IF,c2,b}=102.2232$ meV, the level $\Sigma_{IF,c1}=15.4212$ meV, and the terminal-state level $\Sigma_{A,c}=0.539$ meV corresponding to Fig. A.2(a). In Figs. A.3(a) and A.3(b), the spatial distributions of $\Sigma_{IF,c2,a}$ and $\Sigma_{IF,c2,b}$ are nearly indistinguishable; furthermore, their amplitudes inside the central 9-AGNR segment are extremely weak, indicating that transport through $\Sigma_{IF,c2}$ is strongly suppressed. In contrast, $\Sigma_{IF,c1}$ exhibits a pronounced probability density within the central 9-AGNR segment, making it a more favorable transport channel. As shown in Fig. A.3(c), $\Sigma_{IF,c1}$ effectively behaves as a set of four topological dots (TDs). Figure A.3(d) demonstrates that the state $\Sigma_{A,c}$ is primarily associated with the terminal states of the 27-9-27 AGNRH structure.